| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  73  
 | 
Sat May  8 10:51:41 2021 | 
Akash, Nic, Jelena, Zsolt | Calibration | ESR settings seem different | Settings don't look the same...so we are not 
yet writing files unless sure. | 
    | 
| 
  72  
 | 
Sat May  8 08:44:32 2021 | 
Akash, Nic, Jelena, Zsolt | Accelerator | No Beam! | 08.43am: Beam stopped. HKR looking into it. 
Reason: they mentioned they don't know yet!
 
 | 
   | 
| 
  71  
 | 
Sat May  8 08:16:08 2021 | 
Akash, Nic, Jelena, Zsolt | Analysis | Night shift analysis | (Jelena has analysed these data files)
 
 
E143_245MHz_2021.05.08.00* (-s 5.0 -t 0.5) | 
 16x   | 
| 
  70  
 | 
Sat May  8 08:12:02 2021 | 
Akash, Nic, Jelena, Zsolt | General | SIS, ESR time structure changes | The structure of SIS extraction seems to have 
changed a little...also for ESR injection.
 
 | 
    | 
| 
  69  
 | 
Sat May  8 07:34:08 2021 | 
Akash, Nic, Jelena, Zsolt | General | Zoom has its own mind | Zoom kicked us all out... we managed to get 
back in...screen sharing is back on. | 
   | 
| 
  68  
 | 
Sat May  8 07:14:00 2021 | 
Akash, Nic, Jelena, Zsolt | General | Morning shift entry | Things look ok. No settings have been changed...
 
 
Analysis attempts are going on... | 
   | 
| 
  67  
 | 
Sat May  8 03:08:57 2021 | 
Rui-Jiu Chen | DAQ | Change reference level to -50dBm | Start a new run. The data of new run is in 
:
 
SC_2021-05-07_17-39-16
 | 
   | 
| 
  66  
 | 
Sat May  8 03:05:42 2021 | 
Rui-Jiu Chen | DAQ | Change reference level to -60dBm | Start a new run. The data of new run is in 
:
 
SC_2021-05-07_17-33-29
 | 
   | 
| 
  65  
 | 
Sat May  8 01:43:23 2021 | 
Helmut Weick | Accelerator | 70Se setting | If gamma_t = 1.3955 Tm stays, we will need 
Brho = 6.21902 Tm for 70Se.
 
The corresponding ESR Brho setting will be | 
   | 
| 
  64  
 | 
Sat May  8 01:33:25 2021 | 
Helmut Weick | Simulations | frequency list for 70Se setting | EXCEL table with all possible fragments injected 
into ESR.
 
The path length in ESR, gamma_t and harmonic | 
     | 
| 
  63  
 | 
Sat May  8 01:06:19 2021 | 
Philipp, Marta | General | Screenshot | The signal of the transformer on the lower 
right panel is not really beam related. It 
is mostly noise. Changes are not so expected | 
    | 
| 
  62  
 | 
Sat May  8 00:38:55 2021 | 
Philipp, Marta | General | Screenshot |   | 
    | 
| 
  61  
 | 
Fri May  7 23:08:44 2021 | 
Michael, Kanika | Accelerator | Beam issues for short duration | The beam was lost for a few minutes from 23:01 
to 23:06. It is back now. | 
   | 
| 
  60  
 | 
Fri May  7 22:48:33 2021 | 
Chris | General | still running... | In anticipation of the imminent shift change, 
a snapshot of how things look. | 
     | 
| 
  59  
 | 
Fri May  7 22:00:36 2021 | 
Chris | Accelerator | brief beam loss | Brief interruption in beam, ~21:42-21:51. | 
   | 
| 
  58  
 | 
Fri May  7 20:34:27 2021 | 
Kanika, Michael, Shahab | Runs | back to 15 sec recording | now we are going back to 15 sec operation. 
Cave M (bio) operation is over | 
   | 
| 
  57  
 | 
Fri May  7 20:28:32 2021 | 
Kanika, Michael, Shahab | Analysis | Some pictures from the 410 detector | some pics. | 
 9x   | 
| 
  56  
 | 
Fri May  7 19:56:10 2021 | 
Kanika, Michael, Shahab | Runs | ZOOM session restart | restart zoom session | 
   | 
| 
  55  
 | 
Fri May  7 18:43:13 2021 | 
Kanika, Michael, Shahab | Runs | cave M experiment needs more spills | our pattern causes beam loss while scanning 
the biological probes in the Cave M bio experiment. 
They loose every other spill due to high | 
    | 
| 
  54  
 | 
Fri May  7 18:08:35 2021 | 
Kanika, Sergey, Yuri, Wolfram, Michael, Shahab | Runs | going back to the original energy | going back to the original energy 441 MeV/u
 
 
last filenames with the previous energy:
 | 
   |